OpenStudy (anonymous):

The maximum distance the earth is from the sun is 152,000,000 Km and the minimum distance is 147,000,000 Km. Write an absolute value equation where d represents the maximum and minimum distance between earth and the sun in millions of kilometers.

OpenStudy (anonymous):

Poorly worded question. If d represents the distance from earth to sun as a function of time, d(t), \[147 \le d(t) \le 152\] in millions of km.

OpenStudy (anonymous):

You forgot about the absolute value part of the question

OpenStudy (anonymous):

absolute value is the distance away from 0 so it would be 147 mill and 152 mill its basically a waste of time. thanks albert

OpenStudy (anonymous):

Since distance is always non-negative, there is no need for an absolute value.

OpenStudy (anonymous):

I say again, this is a poorly worded question.

OpenStudy (anonymous):

\[\left| d-149,500,000 \right|\le2,500,000\]

OpenStudy (anonymous):

oops |d−149.5|≤2.5 didn't seethe millions part

OpenStudy (anonymous):

If it's going to be an equation, you need it to be equal. That means when you solve for d, you will get the two numbers given in the problem.

OpenStudy (anonymous):

Nice work. I never would have seen it that way.